# ALPINE: Unveiling the Planning Capability of Autoregressive Learning in Language Models

### v2

@@ 4 Gradient-based Analysis for Path Finding
In case iii), when $\sum_{j}N_{i,j}>0$ and $\bm{W}^{M}_{(i,k)}$ converges to $\infty$- ∞, then the second term in Eq. (10) converges to zero, and it is smaller than $\sum_{j}N_{i,j,k}>0$. Hence, ${\partial\ell(\mathcal{D})\over\partial\bm{W}^{M}_{(i,k)}}$ is negative when $\bm{W}^{M}_{(i,k)}$ converges to $\infty$- ∞. ∎
The theorem directly leads to a theoretical explanation on how the model learns the adjacency and reachability matrices, as explained below.
(a) True adjacency matrix $\bm{A}^{\text{true}}=\bm{A}^{\text{obs}}(\mathcal{D}_{1})=\bm{A}^{\text{obs}}(\mathcal{D}_{2})=\bm{A}^{\text{obs}}(\mathcal{D}_{3})$
(b) Matrix $\bm{W}^{M}$ under $\mathcal{D}_{1}$
(c) Matrix $\bm{W}^{M}$ under $\mathcal{D}_{2}$
(d) Matrix $\bm{W}^{M}$ under $\mathcal{D}_{3}$
Learning the adjacency matrix.
Let $\mathcal{E}(\mathcal{D})$ denote the set of edges appearing in the training dataset $\mathcal{D}$, which corresponds to the observed adjacency matrix $\bm{A}^{\text{obs}}(\mathcal{D})$. For any $(i,k)\in\mathcal{E}(\mathcal{D})$, $\sum_{j}N_{i,j,k}>0$, and for any $(i^{\prime},k^{\prime})\notin\mathcal{E}(\mathcal{D})$, $\sum_{j}N_{i^{\prime},j,k^{\prime}}=0$. Then from the above theorem, under the gradient descent learning paradigm, $\bm{W}^{M}_{(i^{\prime},k^{\prime})}$ will keep decreasing (since its gradient is always positive), while $\bm{W}^{M}_{(i,k)}$ will not (since its gradient becomes negative when its value is sufficiently negative). This tends to make $\bm{W}^{M}_{(i,k)}$ higher than $\bm{W}^{M}_{(i^{\prime},k^{\prime})}$ after training. Note that these terms are weights when predicting the next node: a higher $\bm{W}^{M}_{(i,k)}$ means that “the edge $(i,k)$ exists”, and a lower $\bm{W}^{M}_{(i^{\prime},k^{\prime})}$ means that “the edge $(i^{\prime},k^{\prime})$ does not exist”. By this way, the Transformer model learns the information about the observed adjacency matrix with weight matrix $\bm{W}^{M}$.
[…]
(a) True reachability matrix $\bm{R}^{\text{true}}$
(b) The observed reachability in $\mathcal{D}_{1}$, $\bm{R}^{\text{obs}}(\mathcal{D}_{1})$
(c) Matrix $\bm{W}^{V}$ under $\mathcal{D}_{1}$
(d) The observed reachability in $\mathcal{D}_{2}$, $\bm{R}^{\text{obs}}(\mathcal{D}_{2})$
(e) Matrix $\bm{W}^{V}$ under $\mathcal{D}_{2}$
(f) The observed reachability in $\mathcal{D}_{3}$, $\bm{R}^{\text{obs}}(\mathcal{D}_{3})$
(g) Matrix $\bm{W}^{V}$ under $\mathcal{D}_{3}$
Learning the reachability matrix.
Similar to the process of learning the adjacency matrix, under the gradient descent learning paradigm, $\bm{W}^{V}_{(j^{\prime},k^{\prime})}$ will keep decreasing when $(j^{\prime},k^{\prime})$ is not an observed reachable pairs in the training dataset $\mathcal{D}$. In other words, there is no path in $\mathcal{D}$ in which $j^{\prime}$ is the target and $k^{\prime}$ is a non-source node on the path. On the other hand, when $(j,k)$ is indeed an observed reachable pair, $\bm{W}^{V}_{(j,k)}$ does not keep decreasing. This tends to make $\bm{W}^{V}_{(j,k)}$ higher than $\bm{W}^{V}_{(j^{\prime},k^{\prime})}$ after the training. By this way, the Transformer model captures the structural information of observed reachability matrix with weight matrix $\bm{W}^{V}$. However, our analysis indicates that the model may not learn non-observed reachability relationship even if all the edges are present in the training data. These non-observed reachable pairs $(j,k)$ encompass several cases, which are summarized in Section 3.
Figure 2 shows the correlation between $\bm{W}^{V}$ and the observed reachabilities under different dataset $\mathcal{D}$’s in the above simple experiment. Figure 2(a) is the real reachability matrix of the graph; Figure 2(b) is the observed reachability matrix in dataset $\mathcal{D}_{1}$, and Figure 2(c) is the $\bm{W}^{V}$ matrix under $\mathcal{D}_{1}$; Figure 2(d) is the observed reachability matrix in dataset $\mathcal{D}_{2}$, and Figure 2(e) is the $\bm{W}^{V}$ matrix under $\mathcal{D}_{2}$; and similarly, Figure 2(f) is the observed reachability matrix in dataset $\mathcal{D}_{3}$, and Figure 2(g) is the $\bm{W}^{V}$ matrix under $\mathcal{D}_{3}$. These illustrations shows all the weight matrices $\bm{W}^{V}$ can satisfactorily learn the structural information of the observed reachabilities present in the training datasets.
However, the Transformer models cannot deduce non-observed reachabilities. In particular, we demonstrate that all three types of non-observed reachable pairs as summarized in Section 3 appear in this example: (i) there is no paths that contain the sub-paths from node $5$ to nodes $8$ and $9$ in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, hence the reachable pairs $(8,5)$ and $(9,5)$ are not learned in these two cases (the corresponding entries $(8,5)$ and $(9,5)$ in Figure 2(c) and Figure 2(d) are dark), even though that from these two datasets the model can learn that $5$ reaches $6$, $6$ reaches $7$, $7$ reaches $8$, and $8$ reaches $9$ separately; (ii) there is a path ”0 9 0 2 3 9” in $\mathcal{D}_{2}$, but the reachable pair $(3,2)$ is not learned (the corresponding entry $(3,2)$ in Figure 2(d) is dark red); and (iii) none of these matrices learn the reachable pairs of $(j,k)$ with $k=0,1,4$, since nodes $0,1,4$ never appear as a non-source node in a path.
Predicting the next node on a path.
From Eq.(9), we know that the probability vector for predicting the next node is given as $\mathbf{softmax}(\bm{W}^{M}_{(u_{n},:)}+\bm{W}^{V}_{(u_{2},:)})$, where $u_{n}$ represents the current node, and $u_{2}$ represents the target node. This provides an intuitive explanation for why $\bm{W}^{M}$ learns the observed adjacency matrix, while $\bm{W}^{V}$ learns the observed reachability matrix. The mechanism utilizes $\bm{W}^{M}$ on the current node $u_{n}$ to provide information about the next node it connects to, whereas $\bm{W}^{V}$ is used on the target node to provide information on which nodes can reach the target node. The softmax operation $\mathbf{softmax}(\bm{W}^{M}_{(u_{n},:)}+\bm{W}^{V}_{(u_{2},:)})$ resembles the procedure in Algorithm 1: it predicts the next node $k$ such that both $\bm{W}^{M}_{(u_{n},k)}$ is high (corresponding to $\bm{A}(u_{n},k)=1$) and $\bm{W}^{V}_{(u_{2},k)}$ is high (corresponding to $\bm{R}(u_{2},k)=1$).
In summary, our theoretical analysis demonstrates that a simplified one-layer, one-head auto-regressive Transformer (with perfect attention) can effectively learn crucial adjacency and reachability information from the training data through gradient descent training. Moreover, it can utilize this learned information to predict the next node akin to the decision-making process of a human algorithm designer in similar scenarios. This suggests that, when confronted with the path-finding or more general planning task with a given goal, the Transformer learns the structural information to associate the next step with both the current step and the goal, enabling it to generate the subsequent task step. Nevertheless, the Transformer’s limitation in learning only the observed reachability matrix, without deducing the complete reachability matrix, hints at potential constraints on the goal-oriented information it can acquire. This limitation may result in the Transformer failing to grasp novel reachability relationships derived from the transitivity of reachability relations. In the next section, we will further verify the effectiveness of the Transformer in the path-finding task through extensive empirical evaluations.
5 Empirical Evaluations: Peeking into a Trained Transformer
We conduct extensive experiments on the path-finding task using the general Transformer architecture as described in Section 2.1. The experiments include tests on the overall accuracy of the Transfomer model for the path-finding task, as well as investigation on how well the model learns the attention and the structural information about the adjacency and reachability matrices. In this section, we present these empirical evaluation results, which show that the results derived from our theoretical analysis in Section 4 can be carried over to the general Transformer architecture.
5.1 Datasets
5.1.1 Graphs
The graph is generated randomly on two parameters: the number of nodes $n$, and the probability of edge $p=0.1$. Given these two parameters, we generate a DAG with $n$ nodes as follows: for any $1\leq i<j\leq n$, there is an edge $(i,j)\in\mathcal{E}$ with probability $p$, and the randomness for different edges are independent.
5.1.2 Training Data and Test Data
[…]
5.2 Accuracy on Test Dataset
We train Transformer models on the the aforementioned training dataset and subsequently evaluate performance of these models using the pairs in the test dataset. For test pair $(s,t)$, The correctness of a models’ output is determined based on its validity in terms of syntax and whether it corresponds to a valid path from $s$ to $t$. In our experiments, we employ Transformer models with an embedding size of $d=120$. We conduct tests using various configurations, ranging from 1-layer and 1-head to 6-layer and 6-head, while considering different graph sizes, with number of nodes $n$ ranging from 100 to 500. The accuracy results on all these tests are presented in Figure 3. From these results, we make the following observation:
  • When comparing the five figures, we note that the accuracy tends to decrease as the number of nodes increases. For $n=100$ and $200$, the accuracy consistently remains above $0.95$. However for $n=300$, the accuracy drops to a range between $0.91$ and $0.95$, and for $n=400$ and $500$, the accuracy further declines to a range between $0.80$ and $0.90$ in most cases.
  • When examining at each row, we observe that the accuracy remains relatively stable even as the number of attention heads increases.
  • Upon examining each column, we observe that when the embedding size is sufficiently large in comparison to the graph size (e.g., $n=100,200$), the accuracy remains relatively stable as the number of layers increases. Conversely, when the embedding size is small comparing to the graph size (e.g., $n=300,400,500$), the accuracy shows a slight improvement as the number of layers increases. Specifically, the accuracy with $L=3,4,5,6$ layers tends to be slightly higher than the accuracy achieved with $L=1$ and $2$ layers.
@@ 5.2 Accuracy on Test Dataset
Figure 3: Accuracy on the test dataset with embedding size $d=120$. Takeaway: when the embedding size is large enough, even a Transformer with 1 layer and 1 head can achieve good accuracy; when the embedding size is not large enough, even a Transformer with 6 layers and 6 heads is not enough to achieve good accuracy.
The above observations suggest that both the embedding size and the number of layers have an impact on the model’s accuracy, as they influence the number of parameters in the model. On the one hand, when the embedding size is sufficiently large compared to the graph size, even 1-layer 1-head models would perform well. This coincides with our theoretical analysis in the previous section, which shows that when the embedding size equals to the graph size, the 1-layer and 1-head structure is enough to predict the next nodes accurately. On the other hand, our empirical results show that when the embedding size is small comparing to the graph size, the Transformer architecture may need more parameters to increase the accuracy, which could be achieved by increasing the number of layers in the model. However, increasing the number of layers does not efficiently increase the accuracy a lot, e.g., the accuracy for Transformer with 3 layers is almost the same as the accuracy for Transformer with 6 layers.
@@
Figure 4: The average attention in 1-layer and 1-head Transformer. Takeaway: the Transformer could let its attention concentrate on the target node, especially when the embedding size is large enough.
5.3 Attention
We now look inside the Transformer models, and try to find more evidence that our theoretical analysis reflects the reality. In our theoretical analysis, we assume that the attention is fixed on the target node. Is it true for the Transformer models learned from read data? The corresponding results are shown in Figure 4. These results are obtained by looking into the attention mechanism of the five 1-layer and 1-head Transformer models, and showing the average (taking on the test dataset) matrix of $\textbf{softmax}\left(\frac{{\bm{Q}\bm{K}^{\top}}}{\sqrt{d_{k}}}\right)$, of which the $n^{th}$ row represents the attention vector for predicting the $(n+1)^{th}$ token.
Note that the second column in these figures represents the attention weights on the second token, which corresponding to the target node in our test data. We can see that, when predicting the next tokens, almost all the attention weights are concentrated on this column, especially for those models with higher accuracy (Figure 4(a) for $n=100$ and Figure 4(b) for $n=200$). This demonstrates that indeed the Transformer model learns the correct attention for the path-finding task, and our assumption on the attention for the theoretical analysis is reasonable.
@@ 5.3 Attention
Figure 5: The first 20 rows and columns of $\bm{W}^{M^{\prime}}$ matrix in 1-layer and 1-head Transformer (the red boxes correspond to $1$’s in the adjacency matrix $\bm{A}$), and the average weight gap between edge terms and non-edge terms in $\bm{W}^{M^{\prime}}$. Takeaway: the adjacency matrix is stored in the FFN layer.
5.4 Adjacency Matrix
Our analysis in Section 4 shows that the observed adjacency matrix and the observed reachability matrix are stored in the feed-forward layer and the attention layer. We now verify this on the general 1-layer and 1-head Transformer model that includes all the Transformer mechanisms, such as attention weights, non-linear transformation, token and position embedding.
Note that in the Transformer layer, the output can be written as  [^3] For simplicity, though the layer normalizations $\text{LN}_{1}$,$\text{LN}_{2}$ and $\text{LN}_{t}$ are used in our experiments, we omit them in the equations.
Also noting that we have verified that the attention is concentrated at the second token, then we let $\bm{X}_{2}=\bm{U}_{(2,:)}\bm{W}_{t}$, representing the token embedding of the target node, and $\bm{X}_{n}=\bm{U}_{(n,:)}\bm{W}_{t}$, representing the token embedding of the current node. Then we know that
It is straightforward that $\bm{X}_{n}\bm{W}_{o}$ contains the information of current node, and $\bm{X}_{2}\bm{W}^{V}\bm{W}_{o}$ contains the information of the target node. As for $\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}+\bm{X}_{n}\right)\bm{W}_{o}$, we choose to use its linear approximation as $\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}+\bm{X}_{n}\right)\bm{W}_{o}\approx\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}\right)\bm{W}_{o}+\text{FFN}\left(\bm{X}_{n}\right)\bm{W}_{o}$. As shown by Table 1 (which takes average over all possible $\bm{X}_{2}$’s and $\bm{X}_{n}$’s), this is a good approximation. Then we can treat $\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}\right)\bm{W}_{o}$ as the information of the target node, and $\text{FFN}\left(\bm{X}_{n}\right)\bm{W}_{o}$ as the information of the current node.
Table 1: Average Cosine Similarity of $\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}+\bm{X}_{n}\right)$ and $\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}\right)$ + $\text{FFN}\left(\bm{X}_{n}\right)$
Graph 100 Nodes 200 Nodes 300 Nodes 400 Nodes 500 Nodes
Average Cosine Similarity 0.926 0.924 0.901 0.870 0.889
In this subsection, we want to verify that the adjacency matrix is stored in the feed-forward layer, i.e., in the sum of two terms $\text{FFN}\left(\bm{X}_{n}\right)\bm{W}_{o}+\bm{X}_{n}\bm{W}_{o}$. Let $\bm{W}^{M^{\prime}}$ be the matrix whose $i$-th row is $\text{FFN}\left({\bm{e}}_{i}^{\top}\bm{W}_{t}\right)\bm{W}_{o}+({\bm{e}}_{i}^{\top}\bm{W}_{t})\bm{W}_{o}$, where ${\bm{e}}_{i}$ represents the one-hot column vector for node $i$ (with dimension $M=|\mathcal{V}|+1$). Note that in the simplified Transformer model of Section 4, $\bm{W}^{M^{\prime}}$ is the same as matrix $\bm{W}^{M}$. The results of $\bm{W}^{M^{\prime}}$ are shown in Figure 5.
As we can see, in Figure 5(a), the $\bm{W}^{M^{\prime}}$ matrix and the adjacency matrix are highly aligned: almost all the large entries in the $\bm{W}^{M^{\prime}}$ matrix correspond to real edges, and almost all real edges correspond to large entries in the $\bm{W}^{M^{\prime}}$ matrix. This high accuracy is because the embedding size $d=120$ is higher than the number of nodes $n=100$. If the embedding size is lower than the graph size (Figures 5(b), 5(c), 5(d) and 5(e)), we inevitably lose some accuracy when approximating the adjacency matrix by the product of matrices with rank smaller than the graph size, let alone the non-linear layers’ influence. Even so, there is still high relevance between $\bm{W}^{M^{\prime}}$ and the adjacency matrix: most real edges correspond to large entries in the $\bm{W}^{M^{\prime}}$ matrix.
In Figure 5(f), we show the gap between the average weight corresponds to edges (i.e., the average of $\bm{W}^{M^{\prime}}_{(i,j)}$’s with $i<j$ and $(i,j)\in\mathcal{E}$) and the average weight corresponds to non-edges (i.e., the average of $\bm{W}^{M^{\prime}}_{(i,j)}$’s with $i<j$ and $(i,j)\notin\mathcal{E}$). We can see that in all these five graphs, their gaps keep increasing until convergence, suggesting that weights between edges and non-edges are more easily separated as the learning process proceeds. Moreover, with lower number of nodes, the gap is higher. We believe this is because that when the embedding size is fixed, one can approximate the adjacency matrix better for smaller graphs.
5.5 Reachability Matrix
In this subsection, we want to verify that the reachability matrix is stored in the attention layer, i.e., in the two terms $\bm{X}_{2}\bm{W}^{V}\bm{W}_{o}+\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}\right)\bm{W}_{o}$. Let $\bm{W}^{V^{\prime}}$ be the matrix whose $i^{th}$ row is $({\bm{e}}_{i}^{\top}\bm{W}_{t})\bm{W}^{V}\bm{W}_{o}+\text{FFN}\left(({\bm{e}}_{i}^{\top}\bm{W}_{t})\bm{W}^{V}\right)\bm{W}_{o}$, where ${\bm{e}}_{i}$ represents the one-hot column vector for node $i$ (with dimension $M=|\mathcal{V}|+1$). Note that in the simplified Transformer model of Section 4, $\bm{W}^{V^{\prime}}$ is the same as matrix $\bm{W}^{V}$.
[…]
To further demonstrate that $\bm{R}^{\text{real}}$ is not learned as good as $\bm{R}^{\text{obs}}$, we divide the source-target node pair $(s,t)$ in the test dataset into four categories:
  • Degree 0: $\bm{R}^{\text{obs}}_{(t,s)}=1$.
  • Degree 1: $(s,t)$ is not of degree 0, while $s$ has at least one out-neighbor node $u$ such that $(u,t)$ is of degree 00, i.e. $\bm{R}^{\text{obs}}_{(t,u)}=1$.
  • Degree 2: $(s,t)$ is not of degree 0 and 1, while $s$ has at least one out-neighbor node $u$ such that $(u,t)$ is of degree 1.
  • Degree 3 or more: the remaining $(s,t)$ pairs in the test dataset.
Roughly speaking, in our analysis, when the Transformer is predicting the next node of $s$ for the source-target pair $(s,t)$, it will add the adjacency vector of $s$ and the reachability vector of $t$, and use the sum as the weight vector of the next node. For $(s,t)$ pairs of degree 0 or 1, we know that there is a node $u$ such that $(s,u)\in\mathcal{E}$ and $\bm{R}^{\text{obs}}_{(t,u)}=1$. Then node $u$ will have a large weight, indicating a high accuracy. On the other hand, for $(s,t)$ pairs of degree 2 or more, there is no node $u$ such that both its corresponding entry in the adjacency vector of $s$ and its corresponding entry in the reachability vector of $t$ are large. In this case, the high-weight entry when predicting the next node is either an adjacent node of $s$ or a recorded node that can reach $t$. Since none of the recorded nodes reachable to $t$ is adjacent to $s$, the accuracy for $(s,t)$ pairs of degree 2 or more should be much lower than those of degree 0 or 1.
To see this, we check the accuracy of the Transformer on the $(s,t)$ pairs of the four different categories. The results are shown in Figure 6 (b)-(f). In this figure, each row of the accuracy matrix is further divided into four sub-rows corresponding to the accuracy of degree-0 pairs, degree-1 pairs, degree-2 pairs, and degree-3 or more pairs respectively (in the graph with 100 nodes, there are no test $(s,t)$ pairs in the degree-3 or more category). From these results, we can see that the accuracy for degree-2 pairs and degree-3 or more pairs is much lower than the two other categories in most cases. It indicates that, even with a larger model (e.g. 6-layer 6-head Transformer), the model has a fundamental difficulty to generate paths for high-degree source-target pairs, namely those pairs who can only be connected by concatenating several path segments in the training dataset. In short, the model fails to learn reachability through the transitivity among the reachability relations. This result demonstrates the validity of our theoretical analysis, which shows that the Transformer can only learn observed reachability, and will miss those unobserved reachability deduced from the transitivity of the reachability relation.
5.6 Summary of the Empirical Results
In summary, our extensive empirical evaluation leads to the following conclusions about the auto-regressive Transformer model in achieving the path-finding task: (a) With large enough embedding size, the model can achieve high accuracy in general; (b) The model achieves its performance by concentrating attentions on the target nodes as intended, and learning the information on adjacency and reachability matrices, just as what a human would do and as predicted by our theoretical analysis; and (c) The model may have limitations and fail to learn high-order reachability relations through transitivity, and thus fail to generate paths derived from high-order reachability, as human would expect.
(a) Average Weights in $\bm{W}^{V^{\prime}}$
(b) 100 Nodes
(c) 200 Nodes
(d) 300 Nodes
(e) 400 Nodes
(f) 500 Nodes
Figure 6: The average weights in $\bm{W}^{V^{\prime}}$, and the accuracy for $(s,t)$’s with different degree. Takeaway: the Transformer can only store the observed reachabilities and cannot learn the non-observed reachabilities well.
6 Path-planning in Blocksworld

### v3

@@ Appendix C Proof of Theorem 3
In case iii), when $\sum_{j}N_{i,j}>0$ and $\bm{W}^{M}_{(i,k)}$ converges to $\infty$- ∞, then the second term in Eq. (7) converges to 0, and it is smaller than $\sum_{j}N_{i,j,k}>0$. Hence, ${\partial\ell(\mathcal{D})\over\partial\bm{W}^{M}_{(i,k)}}$ is negative when $\bm{W}^{M}_{(i,k)}$ converges to $\infty$- ∞. ∎
Appendix D The Reason of Choosing $W^{M^{\prime}}$ and $W^{V^{\prime}}$
Note that in the Transformer layer, the output can be written as [^1] For simplicity, though the layer normalizations $\text{LN}_{1}$,$\text{LN}_{2}$ and $\text{LN}_{t}$ are used in our experiments, we omit them in the equations in this section.
Also noting that we have verified that the attention is concentrated at the second token, then we let $\bm{X}_{2}=\bm{U}_{(2,:)}\bm{W}_{t}$, representing the token embedding of the target node, and $\bm{X}_{n}=\bm{U}_{(n,:)}\bm{W}_{t}$, representing the token embedding of the current node.
Then we know that
Table 1: Cosine Similarity of $\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}+\bm{X}_{n}\right)\bm{W}_{o}$ and $\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}\right)\bm{W}_{o}$ + $\text{FFN}\left(\bm{X}_{n}\right)\bm{W}_{o}$
Graph 100 Nodes 200 Nodes 300 Nodes 400 Nodes 500 Nodes
Average Cosine Similarity 0.926 0.924 0.901 0.870 0.889
It is straightforward that $\bm{X}_{n}\bm{W}_{o}$ contains the information of the current node, and $\bm{X}_{2}\bm{W}^{V}\bm{W}_{o}$ contains the information of the target node. As for $\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}+\bm{X}_{n}\right)\bm{W}_{o}$, we choose to use its linear approximation as $\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}+\bm{X}_{n}\right)\bm{W}_{o}\approx\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}\right)\bm{W}_{o}+\text{FFN}\left(\bm{X}_{n}\right)\bm{W}_{o}$. As shown in Table 1 (which takes average over all possible $\bm{X}_{2}$’s and $\bm{X}_{n}$’s), this is a good approximation. Then we can treat $\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}\right)\bm{W}_{o}$ as the information of the target node, and $\text{FFN}\left(\bm{X}_{n}\right)\bm{W}_{o}$ as the information of the current node.
Because of this, we let $\bm{W}^{M^{\prime}}$ be the matrix whose $i$-th row is $\text{FFN}\left({\bm{e}}_{i}^{\top}\bm{W}_{t}\right)\bm{W}_{o}+({\bm{e}}_{i}^{\top}\bm{W}_{t})\bm{W}_{o}$, where ${\bm{e}}_{i}$ represents the one-hot column vector for node $i$ (with dimension $M=|\mathcal{V}|+1$). Note that in the simplified Transformer model of Theorem 3, there is no $\bm{X}_{n}\bm{W}_{o}$ term, and $\bm{W}^{M^{\prime}}$ is the same as matrix $\bm{W}^{M}$. Similarly, we let $\bm{W}^{V^{\prime}}$ be the matrix whose $i^{th}$ row is $({\bm{e}}_{i}^{\top}\bm{W}_{t})\bm{W}^{V}\bm{W}_{o}+\text{FFN}\left(({\bm{e}}_{i}^{\top}\bm{W}_{t})\bm{W}^{V}\right)\bm{W}_{o}$, where ${\bm{e}}_{i}$ represents the one-hot column vector for node $i$ (with dimension $M=|\mathcal{V}|+1$). Note that in the simplified Transformer model of Theorem 3, there is no $\text{FFN}\left(\bm{X}_{2}\bm{W}^{V}\right)\bm{W}_{o}$ term, and $\bm{W}^{V^{\prime}}$ is the same as matrix $\bm{W}^{V}$.
Appendix E Complete Experimental Results on the Synthetic Datasets
@@
Figure 7: Accuracy on the test datasets with embedding size $d=120$.
@@
Figure 8: The average attention in 1-layer and 1-head Transformers.
@@
Figure 9: The first 20 rows and columns of $\bm{W}^{M^{\prime}}$ matrix in the 1-layer and 1-head Transformers (the red boxes correspond to $1$’s in the adjacency matrix $\bm{A}$).
(a) 100 Nodes
(b) 200 Nodes
(c) 300 Nodes
(d) 400 Nodes
(e) 500 Nodes
Figure 10: The accuracy for $(s,t)$’s with different degrees.
In this section, we state the complete experimental results on the synthetic datasets. All these results are conducted on a single A100 GPU.
  • The accuracy results on all these tests are presented in Figure 7.
  • The attention results on all the 1-layer and 1-head Transformers are presented in Figure 8.
  • The results of $\bm{W}^{M^{\prime}}$’s are shown in Figure 9.
  • The accuracy of the Transformers on the $(s,t)$ pairs of the four different categories are shown in Figure 10.
As we can see, all these results are consistent with our conclusions.
Appendix F Path-planning in Blocksworld